OpenStudy (anonymous):

need help in steps. underline each term in the following polynomials and write the number , variable, or expression that each term has in common in the space provided. 5x(3x+1)+7(3x+1) _____________

OpenStudy (ageta):

Let's simplify step-by-step. 5x(3x+1)+7(3x+1) Distribute: =(5x)(3x)+(5x)(1)+(7)(3x)+(7)(1) =15x2+5x+21x+7 Combine Like Terms: =15x2+5x+21x+7 =(15x2)+(5x+21x)+(7) =15x2+26x+7 Answer: =15x2+26x+7

OpenStudy (anonymous):

i dont need to solve it. it says. Step 1 : underline each term in the following polynomials and write the number , variable, or expression that each term has in common in the space provided. A) 5x(3x+1)+7(3x+1) ____________________ Step 2 : Using your answer from step 1 , factor out the number , variable , or expression that each term has in common from the polynomial and multiply it by the fully factored polynomial . show your work